# An algebraic criterion for the vanishing of bounded cohomology

### v4

@@ 3. Wreath products
In this section, we extend the vanishing result of the bounded cohomology of wreath products [Mon22] by considering permutational wreath products. That this is possible was already noticed by Monod [Mon22, Section 4.2]. Though at first glance a minor generalisation, this paves the way to our new powerful criterion for $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity.
@@ 3.1. Permutational wreath products
  • Let $(Y_{0},\mu_{0})$ be the standard Borel probability $\Gamma$-space obtained by endowing $\Gamma$ with a distribution of full support;
@@
[…]
@@ 3.2. Vanishing modulus with coefficients
$\{\Lambda_{i}=\Gamma_{i}\wr_{X_{i}}A_{i}/N_{i}\}_{i\in\mathbb{N}},$
where each $\Gamma_{i}\wr_{X_{i}}A_{i}$ are as in Setup 3.1, and $N_{i}$ is a normal amenable subgroup; and for each $i\in\mathbb{N}$ a separable dual Banach $\Lambda_{i}$-module $E_{i}$ such that
@@
[…]
@@ 4.1. Commuting cyclic conjugates and $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity
A group $\Gamma$ has commuting cyclic conjugates if for every finitely generated subgroup $H\leq\Gamma$ there exist $t\in\Gamma$ and $n\in\mathbb{Z}_{\geq 2}\cup\{\infty\}$ such that $[H,{}^{t^{p}}H]=1$ for $1\leq p<n$ and $[H,t^{n}]=1$. Here we read that $t^{\infty}=1$ and ${}^{t^{p}}H$ denotes the conjugation $t^{p}Ht^{-p}$.
@@
A group has commuting $\mathbb{Z}$-conjugates if for every finitely generated subgroup $H\leq\Gamma$ there exists $t\in\Gamma$ such that $[H,{}^{t^{p}}H]=1$ for all integers $p\geq 1$.
@@
In the definitions above, if the order of $t$ is less than or equal to $n-1$ (or finite when $n=\infty$), then the group $H$ has to be abelian.
@@
Suppose that a group $\Gamma$ has commuting cyclic conjugates. Then the following holds:
@@
    $\Gamma$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic.
@@
    For every finitely generated subgroup $H\leq\Gamma$ there exist a countable amenable group $A$, an infinite transitive $A$-set $X$ and a homomorphism
    $f\colon H\wr_{X}A\to\Gamma$
    such that $f$ restricts to the inclusion on some factor $H$, and $f$ has amenable kernel.
@@
We are now ready to apply Proposition 3.5. Indeed, $H$ is finitely generated and thus countable, $X$ is countably infinite and the amenable group $A$ acts on $X$ with a (unique) infinite orbit. Hence, for every integer $n\geq 1$ there exists a constant $K_{n}\in[0,\infty)$ which does not depend on $H,A,X$ or even $\Gamma$, such that $\|H^{n}_{b}(\operatorname{im}(f);E)\|\leq K_{n}$ for every separable dual Banach $\operatorname{im}(f)$-module $E$. By Proposition 2.24, we conclude that $H^{n}_{b}(\Gamma;E)=0$. ∎
@@
  • $[\Lambda_{i-1},{}^{t_{i}^{p}}\Lambda_{i-1}]=1$ for all $1\leq p<n_{i}$; and
  • $[\Lambda_{i-1},t_{i}^{n_{i}}]=1$.
@@
where the action of $\mathbb{Z}$ on $\mathbb{Z}/n_{i+1}$ is by left translation. By construction, each $A_{i}$ is amenable being an extension of amenable groups. Let $A$ be the direct union of the $A_{i}$. Hence, the group $A$ is countable and amenable, since it is a direct union of countable amenable groups.
@@
To see that the subgroup $B<A$ has infinite index, consider the $A$-conjugates of $A_{1}$. The number of $A_{i}$-conjugates of $A_{i-1}$ is exactly $n_{i}$, so it follows that the number of $A_{i}$-conjugates of $A_{1}$ is at least $n_{2}\cdots n_{i}$. Since by definition of commuting cyclic conjugates each $n_{i}\geq 2$ (Definition 4.1), there are infinitely many $A$-conjugates of $A_{1}$. On the other hand, let us show that $B$ centralizes $A_{1}$. This is clear for $B_{1}$. Recall that an element of $B_{i+1}$ is of the form $((a_{p})_{p\in\mathbb{Z}/n_{i+1}},m)$ where $a_{0}\in B_{i}$ and $m\in n_{i+1}\mathbb{Z}$. Then $a_{0}$ centralizes $A_{1}$ by induction hypothesis, and $a_{p}$ with $p\neq 0$ as well as $m$ centralize $A_{1}$ by definition of the permutational wreath product. Thus all of $B_{i+1}$ centralizes $A_{1}$ and so also their direct union $B$ centralizes $A_{1}$. In conclusion we have proved that the conjugation action of $A$ on the set of its subgroups has the following properties:
@@
This is well-defined since $\langle t_{i+1}^{n_{i+1}}\rangle$ centralizes $\Lambda_{i}$ and therefore $f(A_{i})$ (since $f(A_{i})\leq\Lambda_{i}$ by induction), and $[f(A_{i}),{}^{t_{i+1}^{p}}f(A_{i})]\leq[\Lambda_{i},{}^{t_{i+1}^{p}}\Lambda_{i}]=1$ for all $1\leq p<n_{i+1}$. It is easy to check that it is a homomorphism.
@@ 2. $(ii)$
Let us prove that property (i) is also satisfied. We proceed again by induction, by showing that if $a\in A_{i}$ is such that $[H,{}^{f(a)}H]\neq 1$, then $a\in B_{i}$. For $i=1$, this is by definition of $t_{1}$. Now suppose that $a\in A_{i+1}$ is an element such that $[H,{}^{f(a)}H]\neq 1$. Using the definition of $A_{i+1}=A_{i}\wr_{\mathbb{Z}/n_{i+1}}\mathbb{Z}$, we write $a=((a_{p})_{p\in\mathbb{Z}/n_{i+1}},m)$. Now for $1\leq p<n_{i+1}$ the image of $a_{p}$ lies in ${}^{t_{i+1}^{p}}f(A_{i})\leq{}^{t_{i+1}^{p}}\Lambda_{i}$, which by choice of $t_{i+1}$ commutes with $\Lambda_{i}$, and therefore with $H$. In other words, conjugating by these coordinates has no effect on $H$, and therefore ${}^{f(a)^{-1}}H={}^{t_{i+1}^{-m}}({}^{f(a_{0})^{-1}}H)$. Suppose first that $m\notin n_{i+1}\mathbb{Z}$. Then (${}^{t_{i+1}^{-m}}({}^{f(a_{0})^{-1}}H)\leq{}^{t_{i+1}^{-m}}\Lambda_{i}$, which commutes with $\Lambda_{i}$ and thus with $H$. A standard manipulation shows then that ${}^{f(a)}H$ also commutes with $H$. Since we are assuming that $[H,{}^{f(a)}H]\neq 1$, we get a contradiction. Therefore we must have $m\in n_{i+1}\mathbb{Z}$, which implies (${}^{t_{i+1}^{-m}}({}^{f(a_{0})^{-1}}H)={}^{f(a_{0})^{-1}}H$, since $t_{i+1}^{n_{i+1}}$ centralizes $\Lambda_{i}$. Since we are assuming $[H,{}^{f(a)}H]\neq 1$, or equivalently $[H,{}^{f(a)^{-1}}H]\neq 1$, we obtain $[H,{}^{f(a_{0})^{-1}}H]\neq 1$ and so by induction hypothesis $a_{0}$ actually lies in $B_{i}$. We have shown that $a=((a_{p})_{p\in\mathbb{Z}/n_{i+1}},m)$ is such that $a_{0}\in B_{i}$ and $m\in n_{i+1}\mathbb{Z}$. This shows that $a\in B_{i+1}$ by definition, and concludes the proof. ∎
@@
Let $H\leq\Gamma$ be a finitely generated subgroup, and suppose that there exists a countable amenable group $A$, a morphism $f\colon A\to\Gamma$ and an infinite index subgroup $B<A$ as in condition (2). Note that the property (2).(i) implies that $[^{f(a_{1})}H,^{f(a_{2})}H]=1$ whenever $a_{1}B\neq a_{2}B$. Define $X:=A/B$, with the left translation action of $A$. By the assumptions of (2), $X$ is a transitive infinite $A$-set. We extend $f$ to a homomorphism $f\colon H\wr_{X}A\to\Gamma$ as follows:
@@
Since $f(B)$ centralizes $H$, the value of ${}^{f(a)}h_{aB}$ is independent of the choice of $a$. Moreover, since all of the ${}^{f(a)}H$ pairwise commute, when $a$ runs over a system of representatives, the product is well-defined independently of the order of multiplication. It is easy to check that $f$ is indeed a homomorphism, and by construction it restricts to the inclusion of $H$ on the factor $H_{1B}$. This proves the first part of the claim.
@@
Suppose that $\Gamma$ is a group satisfying condition (3). Let $H\leq\Gamma$ be a finitely generated group and let $f\colon H\wr_{X}A\to\Gamma$ be as required by condition (3). Let $x_{0}\in X$ be such that $f$ restricts to the identity on the $x_{0}$-coordinate. We set $t\coloneqq f(a)$ for some $a\in A$ such that $a\cdot x_{0}\neq x_{0}$. Let $n\geq 2$ be the minimal integer such that $a^{n}\cdot x_{0}=x_{0}$ (possibly $n=\infty$). Then for all $1\leq p<n$ we have $[H,{}^{t^{p}}H]=[f(H_{x_{0}}),{}^{f(a^{p})}f(H_{x_{0}})]=f([H_{x_{0}},H_{a^{p}\cdot x_{0}}])=1$, because $a^{p}\cdot x_{0}\neq x_{0}$; and $[H,t^{n}]=f([H_{x_{0}},a^{n}])=1$ since $a^{n}\cdot x_{0}=x_{0}$. Hence, $\Gamma$ has commuting cyclic conjugates. ∎
@@ 4.3. Constructions
A classical question in bounded cohomology with real coefficients is whether surjective group homomorphisms induce injective maps in bounded cohomology [Bou01, CLM22, FFLM21, FFLM23]. Since in this paper we are working with $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic groups, a weaker formulation of the previous problem is whether quotients of $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic groups are $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic. For this reason, now that we provided a new criterion for bounded acyclicity, it is interesting to note that it does pass to quotients. This shows that quotients of groups with commuting cyclic conjugates are $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic:
@@
Let $\Gamma$ be a group, and let $\pi\colon\Gamma\to\overline{\Gamma}$ be a quotient. For a finitely generated subgroup $\overline{H}\leq\overline{\Gamma}$, choose a preimage for each of the generators to obtain a finitely generated subgroup $H\leq\Gamma$ such that $\pi(H)=\overline{H}$. Since $\Gamma$ has commuting cyclic conjugates, there exist $t\in\Gamma$ and $n\in\mathbb{Z}_{\geq 2}\cup\{\infty\}$ such that $[H,{}^{t^{p}}H]=1$ for $1\leq p<n$ and $[H,t^{n}]=1$. Then by definition of homomorphism we have
$[\overline{H},{}^{\pi(t)^{p}}\overline{H}]=\pi([H,{}^{t^{p}}H])=1$
@@
Using this fact, one can easily construct groups with commuting conjugates but without commuting cyclic conjugates [CFFLM24, Proposition 3.6].
@@
Suppose that the $\Gamma_{i}$ have commuting $\mathbb{Z}$-conjugates. Let $\Gamma\coloneqq\prod_{i\in I}\Gamma_{i}$, and let $H\leq\Gamma$ be finitely generated. Let $H_{i}\leq\Gamma_{i}$ be the projection of $H$ on $\Gamma_{i}$, which is finitely generated. Since $\Gamma_{i}$ has commuting $\mathbb{Z}$-conjugates, there exists $t_{i}\in\Gamma_{i}$ such that $[H_{i},{}^{t_{i}^{p}}H_{i}]=1$ for all $p\geq 1$. Then $t=(t_{i})_{i\in I}\in\Gamma$ satisfies $[H,{}^{t^{p}}H]=1$ for all $p\geq 1$.
@@
Note that the proof does not work for groups with commuting cyclic conjugates, as the numbers $n$ appearing in the definition may be different for different factors. In fact, the product of two groups with commuting cyclic conjugates need not have commuting cyclic conjugates [CFFLM24, Proposition 3.6].
@@
It suffices to show that $[\Gamma,\Gamma]$ has commuting cyclic conjugates, the statement then follows by induction. Let $H\leq[\Gamma,\Gamma]$ be a finitely generated subgroup. Then there exists $t\in\Gamma$ and $n\in\mathbb{Z}_{\geq 2}\cup\{\infty\}$ such that $[H,{}^{t^{p}}H]=1$ for $1\leq p<n$ and $[H,t^{n}]=1$. Now the group $K\coloneqq\langle H,t\rangle$ is also finitely generated, so there exists $s\in\Gamma$ such that $[K,{}^{s}K]=1$. Let $c\coloneqq[t,s]=t\cdot{}^{s}t^{-1}$. Then, for all $g\in K$, it holds ${}^{c}g={}^{t}({}^{{}^{s}t^{-1}}g)={}^{t}g$, and so by induction ${}^{c^{p}}g={}^{t^{p}}g$ for all $p\geq 1$. It follows that $[H,{}^{c^{p}}H]=[H,{}^{t^{p}}H]=1$ for all $1\leq p<n$, and $[H,c^{n}]=[H,t^{n}]=1$. Since $c\in[\Gamma,\Gamma]$, we conclude. ∎

### v5

@@ 3. Wreath products
In this section, we extend the vanishing result of the bounded cohomology of wreath products [Mon22] by considering permutational wreath products. That this is possible was already noticed by Monod [Mon22, Section 4.2]. Though at first glance a minor generalisation, this paves the way to our new criterion for $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity.
@@ 3.1. Permutational wreath products
  • Let $(Y_{0},\mu_{0})$ be the standard Borel probability $\Gamma$-space obtained by endowing $\Gamma$ with a distribution of full support (so $Y_{0}=\Gamma$ as a set);
@@
Proposition 3.2 ($A$ amenable $\Rightarrow$ $Y$ is $\operatorname{\textup{HEZA}}$).
In the Setup 3.1, the $\Gamma\wr_{X}A$-space $Y$ is a $\operatorname{\textup{HEZA}}$-space.
@@ 3.2. Vanishing modulus with coefficients
$\{\Lambda_{i}=\Gamma_{i}\wr_{X_{i}}A_{i}/\penalty 50N_{i}\}_{i\in\mathbb{N}},$
where each $\Gamma_{i}\wr_{X_{i}}A_{i}$ is as in Setup 3.1, and $N_{i}$ is a normal amenable subgroup; and for each $i\in\mathbb{N}$ a separable dual Banach $\Lambda_{i}$-module $E_{i}$ such that
@@
We are going to show that there exists a $\operatorname{\textup{HEZA}}$-space for $G$. Let $Y$ be $G$ itself endowed with a distribution of full support. This is a Zimmer-amenable space for $G$, and for each subgroup $\Gamma_{i}$, by Proposition 2.10.(1). Hence for each $i\in\mathbb{N}$ we can construct a generalised Bernoulli shift $A_{i}\curvearrowright Y^{X_{i}}$ and obtain an action of $\Gamma_{i}\wr_{X_{i}}A_{i}$ on $Y^{X_{i}}$ as explained in Setup 3.1. Proposition 3.2 still applies in this case (we only changed our space $Y$ in the Setup 3.1) and so we have that for each $i\in\mathbb{N}$ the standard Borel probability space $Y^{X_{i}}$ is a $\operatorname{\textup{HEZA}}$-space for $\Gamma_{i}\wr_{X_{i}}A_{i}$.
@@ 4.1. Commuting cyclic conjugates and $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity
A group $\Gamma$ has commuting cyclic conjugates if for every finitely generated subgroup $H\leq\Gamma$ there exist $t\in\Gamma$ and $n\in\mathbb{Z}_{\geq 2}\cup\{\infty\}$ such that $H,{}^{t^{p}}H]=1$ for $1\leq p<n$ and $H,t^{n}]=1$. Here we read that $t^{\infty}=1$ and $^{t^{p}}H$ denotes the conjugation $t^{p}Ht^{-p}$.
@@
A group has commuting $\mathbb{Z}$-conjugates if for every finitely generated subgroup $H\leq\Gamma$ there exists $t\in\Gamma$ such that $H,{}^{t^{p}}H]=1$ for all integers $p\geq 1$.
@@
In the definitions above, if the order of $t$ is less than or equal to n1n-1 (or finite when $n=\infty$), then the group $H$ has to be abelian.
@@
The following holds:
@@
    If $\Gamma$ is a group with commuting cyclic conjugates, then it is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic.
@@
    For every finitely generated subgroup $H\leq\Gamma$ there exist a countable amenable group $A$, an infinite transitive $A$-set $X$ and a homomorphism f:H _X A Γ such that $f$ restricts to the inclusion on some factor $H$, and $f$ has amenable kernel.
@@
We are now ready to apply Proposition 3.5. Indeed, $H$ is finitely generated and thus countable, $X$ is countably infinite and the amenable group $A$ acts on $X$ with a (unique) infinite orbit. Hence, for every integer $n\geq 1$ there exists a constant $K_{n}\in[0,\infty)$ which does not depend on $H,A,X$ or even $\Gamma$, such that $\|H^{n}_{b}(\operatorname{im}(f);E)\|\leq K_{n}$ for every separable dual Banach $\operatorname{im}(f)$-module $E$.
Every finitely generated $H\leq\Gamma$ is contained in a group of the form $\operatorname{im}(f)$ as above. Hence, applying Proposition 2.24 with $\Gamma_{i}=\operatorname{im}(f)$, we conclude that $H^{n}_{b}(\Gamma;E)=0$. ∎
@@
  • $\Lambda_{i-1},{}^{t_{i}^{p}}\Lambda_{i-1}]=1$ for all $1\leq p<n_{i}$; and
  • $\Lambda_{i-1},t_{i}^{n_{i}}]=1$.
@@
where the action of $\mathbb{Z}$ on $\mathbb{Z}/n_{i+1}$ is by left translation. By construction, each $A_{i}$ is amenable being an extension of amenable groups. Consider the inclusion of $A_{i}$ into $A_{i+1}$, concentrated at the coordinate $0\in\mathbb{Z}/n_{i+1}$. Let $A$ be the direct union of the $A_{i}$. Hence, the group $A$ is countable and amenable, since it is a direct union of countable amenable groups.
@@
To see that the subgroup $B<A$ has infinite index, consider the $A$-conjugates of $A_{1}$. First note that there are exactly $n_{i}$ conjugates of $A_{i-1}$ inside $A_{i}$, which have pairwise trivial intersection: these are the copies of $A_{i-1}$ concentrated at the different coordinates of $\mathbb{Z}/n_{i}$. It follows that the number of $A_{i}$-conjugates of $A_{1}$ is at least $n_{2}\cdots n_{i}$. Since by definition of commuting cyclic conjugates each $n_{i}\geq 2$ (Definition 4.1), there are infinitely many $A$-conjugates of $A_{1}$. On the other hand, let us show that $B$ centralizes $A_{1}$. This is clear for $B_{1}$. Recall that an element of $B_{i+1}$ is of the form $((a_{p})_{p\in\mathbb{Z}/n_{i+1}},m)$ where $a_{0}\in B_{i}$ and $m\in n_{i+1}\mathbb{Z}$. Then $a_{0}$ centralizes $A_{1}$ by induction hypothesis, and $a_{p}$ with $p\neq 0$ as well as $m$ centralize $A_{1}$ by definition of the permutational wreath product. Thus all of $B_{i+1}$ centralizes $A_{1}$ and so also their direct union $B$ centralizes $A_{1}$. In conclusion we have proved that the conjugation action of $A$ on the set of its subgroups has the following properties:
@@
Let us check that this is a homomorphism. By induction it is a homomorphism on each summand $A_{i}$. Recall that $\langle t_{i+1}^{n_{i+1}}\rangle$ centralizes $\Lambda_{i}$ and therefore $f(A_{i})$ (since $f(A_{i})\leq\Lambda_{i}$ by induction). Recall also that $f(A_{i}),{}^{t_{i+1}^{p}}f(A_{i})]\leq[\Lambda_{i},{}^{t_{i+1}^{p}}\Lambda_{i}]=1$ for all $1\leq p<n_{i+1}$. Applying a suitable conjugation, we see that the subgroups $^{t_{i+1}^{p}}f(A_{i})$ pairwise commute, hence $f$ is a homomorphism on the direct sum of the $A_{i}$. Finally, the action of $\mathbb{Z}$ by conjugacy on the summands is respected, by construction and because $t_{i+1}^{n_{i+1}}$ centralizes each summand.
@@ 2. $(ii)$
[…]
@@
Let $H\leq\Gamma$ be a finitely generated subgroup, and suppose that there exists a countable amenable group $A$, a morphism $f\colon A\to\Gamma$ and an infinite index subgroup $B<A$ as in condition (2). Note that the property (2).(i) implies that $^{f(a_{1})}H,^{f(a_{2})}H]=1$ whenever $a_{1}B\neq a_{2}B$. Define $X:=A/B$, with the left translation action of $A$. By the assumptions of (2), $X$ is a transitive infinite $A$-set. We extend $f$ to a homomorphism $f\colon H\wr_{X}A\to\Gamma$ as follows:
@@
Since $f(B)$ centralizes $H$, the value of $^{f(a)}h_{aB}$ is independent of the choice of $a$. Moreover, since all of the $^{f(a)}H$ pairwise commute, when $a$ runs over a system of representatives, the product is well-defined independently of the order of multiplication. Finally, the action by conjugacy of $A$ on the $H$ summands corresponds equivariantly to the action by conjugacy of $f(A)$ on the commuting groups $^{f(a)}H:aB\in X$. This implies that $f$ is indeed a homomorphism, and by construction it restricts to the inclusion of $H$ on the factor $H_{1B}$. This proves the first part of the claim.
@@
Suppose that $\Gamma$ is a group satisfying condition (3). Let $H\leq\Gamma$ be a finitely generated group and let $f\colon H\wr_{X}A\to\Gamma$ be as required by condition (3). Let $x_{0}\in X$ be such that $f$ restricts to the identity on the $x_{0}$-coordinate. We set $t\coloneqq f(a)$ for some $a\in A$ such that $a\cdot x_{0}\neq x_{0}$. Let $n\geq 2$ be the minimal integer such that $a^{n}\cdot x_{0}=x_{0}$ (possibly $n=\infty$). Then for all $1\leq p<n$ we have $H,{}^{t^{p}}H]=[f(H_{x_{0}}),{}^{f(a^{p})}f(H_{x_{0}})]=f([H_{x_{0}},H_{a^{p}\cdot x_{0}}])=1$, because $a^{p}\cdot x_{0}\neq x_{0}$; and $H,t^{n}]=f([H_{x_{0}},a^{n}])=1$ since $a^{n}\cdot x_{0}=x_{0}$. Hence, $\Gamma$ has commuting cyclic conjugates. ∎
@@ 4.3. Constructions
A classical question in bounded cohomology with real coefficients is whether surjective group homomorphisms induce injective maps in bounded cohomology [Bou01, CLM22, FFLM24, FFLM23]. Since in this paper we are working with $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic groups, a weaker formulation of the previous problem is whether quotients of $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic groups are $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic. For this reason, now that we provided a new criterion for bounded acyclicity, it is interesting to note that it does pass to quotients. This shows that quotients of groups with commuting cyclic conjugates are $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic:
@@
Let $\Gamma$ be a group, and let $\pi\colon\Gamma\to\overline{\Gamma}$ be a quotient. For a finitely generated subgroup $\overline{H}\leq\overline{\Gamma}$, choose a preimage for each of the generators to obtain a finitely generated subgroup $H\leq\Gamma$ such that $\pi(H)=\overline{H}$. Since $\Gamma$ has commuting cyclic conjugates, there exist $t\in\Gamma$ and $n\in\mathbb{Z}_{\geq 2}\cup\{\infty\}$ such that $H,{}^{t^{p}}H]=1$ for $1\leq p<n$ and $H,t^{n}]=1$. Then by definition of homomorphism we have
$\overline{H},{}^{\pi(t)^{p}}\overline{H}]=\pi([H,{}^{t^{p}}H])=1$
@@
Using this fact, one can easily construct groups with commuting conjugates but without commuting cyclic conjugates [CFFLM25, Proposition 3.6].
@@
Suppose that the $\Gamma_{i}$ have commuting $\mathbb{Z}$-conjugates. Let $\Gamma\coloneqq\prod_{i\in I}\Gamma_{i}$, and let $H\leq\Gamma$ be finitely generated. Let $H_{i}\leq\Gamma_{i}$ be the projection of $H$ on $\Gamma_{i}$, which is finitely generated. Since $\Gamma_{i}$ has commuting $\mathbb{Z}$-conjugates, there exists $t_{i}\in\Gamma_{i}$ such that $H_{i},{}^{t_{i}^{p}}H_{i}]=1$ for all $p\geq 1$. Then $t=(t_{i})_{i\in I}\in\Gamma$ satisfies $H,{}^{t^{p}}H]=1$ for all $p\geq 1$.
@@
Note that the proof does not work for groups with commuting cyclic conjugates, as the numbers $n$ appearing in the definition may be different for different factors. In fact, the product of two groups with commuting cyclic conjugates need not have commuting cyclic conjugates [CFFLM25, Proposition 3.6].
@@
It suffices to show that $\Gamma,\Gamma]$ has commuting cyclic conjugates, the statement then follows by induction. Let $H\leq[\Gamma,\Gamma]$ be a finitely generated subgroup. Then there exists $t\in\Gamma$ and $n\in\mathbb{Z}_{\geq 2}\cup\{\infty\}$ such that $H,{}^{t^{p}}H]=1$ for $1\leq p<n$ and $H,t^{n}]=1$. Now the group $K\coloneqq\langle H,t\rangle$ is also finitely generated, so there exists $s\in\Gamma$ such that $K,{}^{s}K]=1$. Let $c\coloneqq[t,s]=t\cdot{}^{s}t^{-1}$. Then, for all $g\in K$, it holds $^{c}g={}^{t}({}^{{}^{s}t^{-1}}g)={}^{t}g$, and so by induction $^{c^{p}}g={}^{t^{p}}g$ for all $p\geq 1$. It follows that $H,{}^{c^{p}}H]=[H,{}^{t^{p}}H]=1$ for all $1\leq p<n$, and $H,c^{n}]=[H,t^{n}]=1$. Since $c\in[\Gamma,\Gamma]$, we conclude. ∎